# Codimension-Two Spiral Spin-Liquid in the Effective Honeycomb-Lattice Compound Cs3Fe2Cl9

### v1

@@ Abstract
A codimension-two spiral spin-liquid is a correlated paramagnetic state with one-dimensional ground state degeneracy hosted within a three-dimensional lattice. Here, via neutron scattering experiments and numerical simulations, we establish the existence of a codimension-two spiral spin-liquid in the effective honeycomb-lattice compound Cs3Fe2Cl9 and demonstrate the selective visibility of the spiral surface through phase tuning. In the long-range ordered regime, competing spiral and spin density wave orders emerge as a function of applied magnetic field, among which a possible order-by-disorder transition is identified.
pacs:
@@
Introduction. A spiral spin-liquid (SSL) is an exotic type of correlated paramagnetic state where the low energy dynamics consist of collective spiral correlations Bergman et al. (2007); Lee and Balents (2008); Mulder et al. (2010); Zhang and Lamas (2013); Niggemann et al. (2020); Attig and Trebst (2017); Balla et al. (2020, 2019); Lee and Balents (2008); Chen (2017); Yao et al. (2021); Huang et al. (2022); Buessen et al. (2018); Liu et al. (2020); Mohylna et al. (2022). A characteristic feature of a SSL is that the propagation vectors of the degenerate spiral ground states form a continuous surface in reciprocal space Bergman et al. (2007). Such an unusual yet clearly defined feature has stimulated a strong interest by the community to experimentally identify and understand SSLs in real materials Tristan et al. (2005); T. Suzuki and H. Nagai and M. Nohara and H. Takagi (2007); Krimmel et al. (2009); Matsuda et al. (2010); MacDougall et al. (2011); Zaharko et al. (2011); Nair et al. (2014); MacDougall et al. (2016); Ge et al. (2017); Graham et al. (2023); Gao et al. (2017, 2020); Rosales et al. (2022); Guratinder et al. (2022); Chamorro et al. (2018); Bai et al. (2019); Tsurkan et al. (2021); Haraguchi et al. (2019); Abdeldaim et al. (2020); Otsuka et al. (2020); Wessler et al. (2020); Bordelon et al. (2021); Balz et al. (2016); Pohle et al. (2021); Takahashi et al. (2024); Gao et al. (2022a); Cole et al. (2023); Gao et al. (2022b); Hsieh et al. (2022). Through experimental and theoretical studies, compounds with a bipartite lattice, e.g. the honeycomb Balz et al. (2016); Pohle et al. (2021); Gao et al. (2022a); Cole et al. (2023) and diamond Graham et al. (2023); Balz et al. (2016); Pohle et al. (2021); Gao et al. (2017); Bai et al. (2019); Guratinder et al. (2022) lattices, have been demonstrated as the most fertile hosts of SSLs.
@@
Figure 1: (a) AB-stacked triangular bilayers formed by the Fe3+ ions in Cs3Fe2Cl9. Atoms belonging to the neighboring bilayers are shown in red and blue, respectively. The $J_{1}$, $J_{2}$, and $J_{3}$ bonds are shown by thick black lines, thin black lines, and thin gray lines, respectively. Arrows indicate the spin directions of the collinear ground state with $\bm{q}=(\frac{1}{2},0,0)$. (b) The AB-stacked triangular bilayers viewed along the $c$ axis. (c) Refinement result of the powder neutron diffraction data measured on POWGEN at $T$ = 1.6 K. Data points are shown as red circles. The calculated pattern is shown as the black solid line. The vertical bars indicate the positions of the structural (upper) and magnetic (lower) Bragg peaks for Cs3Fe2Cl9. The blue line at the bottom shows the difference of measured and calculated intensities. The goodness-of-fit parameters are $R_{\textrm{p}}=18.6\%$ and $R_{\textrm{wp}}=10.2\%$.
@@
Identifying novel SSL hosts is crucial for the realization of exotic spin textures like skyrmions Gao et al. (2020); Rosales et al. (2022); Takeda et al. (2024) and subdimensional quasiparticles like fractons Yan and Reuther (2022); Pretko et al. (2020), and will also establish new candidate compounds to study the thermal and quantum order-by-disorder (ObD) transitions that are elusive in real materials Villain et al. (1980); Henley (1989); Green et al. (2018); Bergman et al. (2007); Mulder et al. (2010). According to theoretical studies, SSLs can be classified by their codimension, a quantity that characterizes the dimensional difference between the spiral surface and the host system Yao et al. (2021). Experimentally identified SSLs, including those observed on the diamond and honeycomb lattices, exhibit either a two-dimensional (2D) spiral surface on a three-dimensional (3D) lattice Gao et al. (2017); Bai et al. (2019); Guratinder et al. (2022); Gao et al. (2022b); Graham et al. (2023) or a one-dimensional (1D) spiral surface, i.e. a degenerate line, on a 2D lattice Balz et al. (2016); Pohle et al. (2021); Takahashi et al. (2024); Gao et al. (2022a), thus all falling in the codimension one category. Although codimension-two SSLs have been predicted to exist on the AB-stacked triangular lattice Yao et al. (2021); Liu et al. (2016); Hoang and Diep (2012); Rastelli et al. (1983), their experimental realizations remain an open question.
@@
Spin Dynamics and Modeling. To determine the exchange coupling strengths, inelastic neutron scattering (INS) experiments were performed at CNCS Ehlers et al. (2011) at SNS of ORNL sup . Figure 2(a) presents the INS spectra collected at $T=2$ K with an incident neutron energy of $E_{i}=3.32$ meV. Even for the powder sample, it is clear that there are two highly dispersive magnon modes centered around $E\sim 0.6$ and 1.8 meV energy transfer. From Fig. 2(a) and also the $E_{i}=1.0$ meV data in the Supplemental Material sup , an excitation gap of $\Delta\sim 0.2$ meV is observed at wavevector transfer $Q\sim 0.6$-1, suggesting the existence of a uniaxial single-ion anisotropy (SIA) that stabilizes the collinear ground state.
@@
Using linear spin wave theory as implemented in the SpinW program Toth and Lake (2015), $\chi^{2}$-fits to the INS spectra were performed to analyze the spin interactions. A minimal $J_{123}$-$D_{z}$ model with Hamiltonian $\mathcal{H}=\sum_{\langle ij\rangle\in n}J_{n}\bm{S}_{i}\cdot\bm{S}_{j}+D_{z}(S_{z})^{2}$ is employed in our calculations, which considers Heisenberg exchange interactions up to the third neighbors as shown in Fig. 1 plus a uniaxial SIA term, $D_{z}$. As compared in Fig. 2, the $J_{123}$-$D_{z}$ model with fitted coupling strengths of $J_{1}=-0.238(2)$, $J_{2}=0.085(2)$, $J_{3}=0.053(1)$, and $D_{z}=-0.037(2)$ meV reproduces the INS spectra. The dominant FM $J_{1}$ together with comparable strengths of $J_{2}$ and $J_{3}$ favor SSLs Mulder et al. (2010); Yao et al. (2021); Liu et al. (2016), while the relatively high magnitude of $|D_{z}|\sim 0.65J_{3}$ indicates the importance of the SIA.
Figure 2: (a) Inelastic neutron scattering spectra $S(q,\omega)$ for Cs3Fe2Cl9 powders at $T=2$ K measured on CNCS. (b) Simulated spectra for the $J_{123}$-$D_{z}$ model using linear spin wave theory.
@@
Codimension-two SSLs. Figure 3(a) presents the diffuse neutron scattering pattern for Cs3Fe2Cl9 in the ($h$, $k$, 0) plane. Data were collected from CORELLI Ye et al. (2018) at SNS of ORNL using an $\sim 8$ mg crystal sup . At $T=6$ K, triangular shaped lobes are observed around the $K$-($\frac{1}{3}$, $\frac{1}{3}$, 0) points. As indicated by the dashed curves, the shape of the spiral surface is reproduced by a $J_{1}^{h}$-$J_{2}^{h}$ honeycomb-lattice model with $J_{2}^{h}/J_{1}^{h}=0.62$, where $J_{1}^{h}$ $(J_{2}^{h})$ is equivalent to $J_{2}$ ($J_{3}$) in the $J_{123}$-$D_{z}$ model. The equal-time spin correlations for the fitted $J_{123}$-$D_{z}$ model can be calculated using the self-consistent Gaussian approximation (SCGA) method Canals and Garanin (2001); sup . Since the critical correlations are underestimated in the SCGA method, a reduced $T=5$ K is assumed in the calculations to better describe the experimental data sup . As compared in Fig. 3(a), the calculated and experimental intensity distributions agree well with each other, both following the spiral surface of a SSL on a honeycomb lattice.
@@
[…]
@@
[…]
Due to the limited number of magnetic reflections collected on WAND2, we resort to classical Monte Carlo simulations to understand the rich phases in Cs3Fe2Cl9. As explained in the Supplemental Material sup , the fitted $J_{123}$-$D_{z}$ model exhibits a relatively simple phase diagram that contains only two main phases with $\bm{q}_{\textrm{I}}=(\frac{1}{2},0,0)$. Therefore, starting from the fitted $J_{123}$-$D_{z}$ model, we examined the impacts of weak perturbations from isotropic exchange interactions up to the tenth neighbors one-by-one. The exchange paths of each perturbations and the representative theoretical phase diagrams are presented in the Supplemental Material sup .
Following this method, we found that by adding a weak fifth-neighbor coupling $J_{5}=-0.008$ meV that only marginally impacts the INS spectra and diffuse scattering patterns sup , the calculated phase diagram, shown in Fig. 4(f), reproduces the main phases observed in experiments. Phases VI and VI′′ (Phase V) adjacent to phases I and VII exhibit similar diffraction pattern as that in phase VI (V). These primed phases, V, VI, and VI′′, are absent in experiments and may arise from finite size effects in our simulations. The magnetic orders in the main phases are described in Figs. 4(g)-(j), with their corresponding Ansatz and structure factor plots presented in the Supplemental Material sup . As is consistent with the experimental data, similar propagation vectors are observed in phases III-V. Among them, phases III and IV are determined to be of the spiral-type orders by inspecting the spin configuration snapshots in the Monte Carlo simulations, where the transition between phases III and IV can be attributed to the modulation of the $\bm{q}=0$ magnitude sup . As a contrast, both phases V and VI are revealed to be of the SDW orders with spins along the $c$ axis despite their different $\bm{q}$ vectors. In phases III-V where sufficient numbers of magnetic Bragg peaks can be extracted in our neutron diffraction experiments, the proposed magnetic structures agree with the experimental dataset sup , thus verifying the results of our Monte Carlo simulations.
Since $\bm{q}_{\textrm{VI}}$ in phase VI is the predicted propagation vector of the ObD theory Mulder et al. (2010); Liu et al. (2020), it is tantalizing to ascribe the transition from the SSL state to phase VI as entropy-driven. Compared to the theoretical phase diagram for the fitted $J_{123}$-$D_{z}$ model, it can be concluded that in the presence of a magnetic field, the $J_{5}$ perturbations favor orders with propagation vectors over the spiral surface, among which $\bm{q}_{\textrm{VI}}$ is stabilized by thermal fluctuations at higher temperatures. The proposed SDW character of phase VI, however, poses a challenge in accurately determining the free energy for states above $T_{N}$ under the current theoretical framework Bergman et al. (2007), making it difficult to exclude perturbations other than thermal fluctuations.
Conclusion. Our neutron scattering experiments establish Cs3Fe2Cl9 as a host of codimension-two SSLs, where the high codimension is manifested through the phase tuning of the visibility of the spiral surface. By combining neutron diffraction and classical Monte Carlo simulations, we clarify the rich field-induced phases as competing spiral and SDW orders, revealing Cs3Fe2Cl9 as a candidate compound to study the ObD transition in frustrated magnets.

### v2

@@ Abstract
A codimension-two spiral spin-liquid is a correlated paramagnetic state with one-dimensional ground state degeneracy hosted within a three-dimensional lattice. Here, via neutron scattering experiments and numerical simulations, we establish the existence of a codimension-two spiral spin-liquid in the effective honeycomb-lattice compound Cs3Fe2Cl9, which demonstrates a novel path to spiral spin-liquids by overcoming the long-standing impediment of weak further-neighbor interactions. In the long-range ordered regime, competing spiral and spin density wave orders emerge as a function of applied magnetic field, among which a possible order-by-disorder transition is identified.
@@
Introduction. A spiral spin-liquid (SSL) is an exotic type of correlated paramagnetic state where the low energy dynamics consist of collective spiral correlations Bergman et al. (2007); Lee and Balents (2008); Mulder et al. (2010); Zhang and Lamas (2013); Niggemann et al. (2020); Attig and Trebst (2017); Balla et al. (2020, 2019); Lee and Balents (2008); Chen (2017); Yao et al. (2021); Huang et al. (2022); Buessen et al. (2018); Liu et al. (2020); Mohylna et al. (2022). A characteristic feature of a SSL is that the propagation vectors of the degenerate spiral ground states form a continuous surface in reciprocal space Bergman et al. (2007). Such an unusual yet clearly defined feature has stimulated a strong interest by the community to experimentally identify and understand SSLs in real materials Tristan et al. (2005); T. Suzuki and H. Nagai and M. Nohara and H. Takagi (2007); Krimmel et al. (2009); Matsuda et al. (2010); MacDougall et al. (2011); Zaharko et al. (2011); Nair et al. (2014); MacDougall et al. (2016); Ge et al. (2017); Graham et al. (2023); Gao et al. (2017, 2020); Rosales et al. (2022); Guratinder et al. (2022); Chamorro et al. (2018); Bai et al. (2019); Tsurkan et al. (2021); Haraguchi et al. (2019); Abdeldaim et al. (2020); Otsuka et al. (2020); Wessler et al. (2020); Bordelon et al. (2021); Balz et al. (2016); Pohle et al. (2021); Takahashi et al. (2024); Gao et al. (2022a); Cole et al. (2023); Gao et al. (2022b); Hsieh et al. (2022); Wan et al. (2024); Andriushin et al. (2025). Through experimental and theoretical studies, compounds with a bipartite lattice, e.g. the honeycomb Balz et al. (2016); Pohle et al. (2021); Gao et al. (2022a); Cole et al. (2023) and diamond Graham et al. (2023); Balz et al. (2016); Pohle et al. (2021); Gao et al. (2017); Bai et al. (2019); Guratinder et al. (2022) lattices, have been demonstrated as the most fertile hosts of SSLs, where spiral correlations arise from the competition between the inter- and intra-sublattice interactions.
@@
Figure 1: (a) AB-stacked triangular bilayers formed by the Fe3+ ions in Cs3Fe2Cl9. Atoms belonging to the neighboring bilayers are shown in red and blue, respectively. The $J_{1}$, $J_{2}$, and $J_{3}$ bonds are shown by thick black lines, thin black lines, and thin gray lines, respectively. The $J_{4}$ and $J_{5}$ bonds are indicated by red curved lines. Arrows indicate the spin directions of the collinear ground state with $\bm{q}=(\frac{1}{2},0,0)$. (b) The AB-stacked triangular bilayers viewed along the $c$ axis. (c) Refinement result of the powder neutron diffraction data measured on POWGEN at $T$ = 1.6 K. Data points are shown as red circles. The calculated pattern is shown as the black solid line. The vertical bars indicate the positions of the structural (upper) and magnetic (lower) Bragg peaks for Cs3Fe2Cl9. The blue line at the bottom shows the difference of measured and calculated intensities. The goodness-of-fit parameters are $R_{\textrm{p}}=18.6\%$ and $R_{\textrm{wp}}=10.2\%$.
Figure 2: (a,c,e,g) Inelastic neutron scattering spectra $S(q,\omega)$ for Cs3Fe2Cl9 powders measured on CNCS with $E_{i}=3.32$ meV at $T=2$ (a) and 10 K (c), and with $E_{i}=1.0$ meV at $T=2$ (e) and 10 K (g). (b,f) Simulated spectra for the fitted $J_{1\textrm{-}5}$-$D_{z}$ model using linear spin wave theory (LSWT). The spectra in (b) and (f) are convolved with the instrumental energy resolution for $E_{i}=3.32$ and 1.0 meV, respectively. (d,h) Simulated spectra for the fitted $J_{1\textrm{-}5}$-$D_{z}$ model using the Landau-Langevin-Gilbert (LLG) method calculated at $T=10$ K.
@@
Identifying novel SSL hosts is crucial for the realization of exotic spin textures like skyrmions Gao et al. (2020); Rosales et al. (2022); Takeda et al. (2024) and subdimensional quasiparticles like fractons Yan and Reuther (2022); Pretko et al. (2020), and will also establish new candidate compounds to study the thermal and quantum order-by-disorder (ObD) transitions that are elusive in real materials Villain et al. (1980); Henley (1989); Green et al. (2018); Bergman et al. (2007); Mulder et al. (2010). According to theoretical studies, SSLs can be classified by their codimension, a quantity that characterizes the dimensional difference between the spiral surface and the host system Yao et al. (2021). Experimentally identified SSLs, including those observed on the diamond and honeycomb lattices, exhibit either a two-dimensional (2D) spiral surface on a three-dimensional (3D) lattice Gao et al. (2017); Bai et al. (2019); Guratinder et al. (2022); Gao et al. (2022b); Graham et al. (2023) or a one-dimensional (1D) spiral surface, i.e. a degenerate line, on a 2D lattice Balz et al. (2016); Pohle et al. (2021); Takahashi et al. (2024); Gao et al. (2022a), thus all falling in the codimension one category. Although codimension-two SSLs have been predicted to exist on the AB-stacked triangular lattice Yao et al. (2021); Liu et al. (2016); Hoang and Diep (2012); Rastelli et al. (1983), their experimental realizations remain an open question. If proven true, this novel approach to spiral spin-liquids, where the intra-sublattice coupling arises from the relatively strong intralayer interactions, will relieve the long-standing obstacle of insufficient frustration on the original honeycomb lattice Gao et al. (2022a).
@@
Spin Dynamics and Modeling. To determine the exchange coupling strengths, inelastic neutron scattering (INS) experiments were performed at CNCS Ehlers et al. (2011) at SNS of ORNL sup . Figure 2(a) presents the INS spectra collected with an incident neutron energy of $E_{i}=3.32$ (a,c) and 1.0 meV (e,g) at temperatures $T=2$ (a,e) and 10 K (c,g), which are below and above $T_{N}$, respectively. Even for the powder sample, it is clear that there are two highly dispersive magnon modes centered around $E\sim 0.6$ and 1.8 meV energy transfer. From Fig. 2(e), an excitation gap of $\Delta\sim 0.2$ meV is observed at wavevector transfer $Q\sim 0.6$ Å-1, suggesting the existence of a uniaxial single-ion anisotropy (SIA) that stabilizes the collinear ground state.
@@
Using linear spin wave theory as implemented in the SpinW program Toth and Lake (2015), $\chi^{2}$-fits to the INS spectra were performed to analyze the spin interactions. As explained in the Supplemental Material sup , a minimal $J_{1\textrm{-}5}$-$D_{z}$ model with Hamiltonian $\mathcal{H}=\sum_{\langle ij\rangle\in n}J_{n}\bm{S}_{i}\cdot\bm{S}_{j}+D_{z}(S_{z})^{2}$ is employed in our calculations, which considers Heisenberg exchange interactions up to the fifth neighbors as shown in Fig. 1 plus a uniaxial SIA term, $D_{z}$. As compared in Fig. 2(a,b,e,f), the $J_{1\textrm{-}5}$-$D_{z}$ model with fitted coupling strengths of $J_{1}=-0.231(2)$, $J_{2}=0.082(2)$, $J_{3}=0.059(1)$, $J_{4}=0.0050(1)$, $J_{5}=0.0013(1)$ and $D_{z}=-0.032(2)$ meV reproduces the INS spectra. The dominant FM $J_{1}$ together with comparable strengths of $J_{2}$ and $J_{3}$ favor SSLs Mulder et al. (2010); Yao et al. (2021); Liu et al. (2016), while the relatively high magnitude of $|D_{z}|\sim 0.54J_{3}$ indicates the importance of the SIA in stabilizing the ground state.
The accuracy of the fitted model is further verified through the comparison of the INS spectra at $T=10$ K shown in Fig. 2(c,d,g,h), where the calculations were performed using the Landau-Lifshitz method as implemented in the JuliaMD program Jul . At elevated temperatures, the magnon mode at higher energy becomes broadened and softened, and the magnon band at lower energy collapses towards the elastic line, which indicates the emergence of strong spin correlation in the paramagnetic regime that can be directly detected through diffuse neutron scattering.
@@
Codimension-two SSLs. Figure 3(a) presents the diffuse neutron scattering pattern for Cs3Fe2Cl9 in the ($h$, $k$, 0) plane. Data were collected from CORELLI Ye et al. (2018) at SNS of ORNL using an $\sim 8$ mg crystal sup . At $T=6$ K, triangular shaped lobes are observed around the KK-($\frac{1}{3}$, $\frac{1}{3}$, 0) points. As indicated by the dashed curves, the shape of the spiral surface is reproduced by a $J_{1}^{h}$-$J_{2}^{h}$ honeycomb-lattice model with $J_{2}^{h}/J_{1}^{h}=0.72$, where $J_{1}^{h}$ $(J_{2}^{h})$ is equivalent to $J_{2}$ ($J_{3}$) in the $J_{1\textrm{-}5}$-$D_{z}$ model. The equal-time spin correlations for the fitted $J_{1\textrm{-}5}$-$D_{z}$ model can be calculated using the self-consistent Gaussian approximation (SCGA) method Canals and Garanin (2001); sup . Since the critical correlations are underestimated in the SCGA method, a reduced $T=5$ K is assumed in the calculations to better describe the experimental data. As compared in Fig. 3(a), the calculated and experimental intensity distributions agree well with each other, both following the spiral surface of a SSL on a honeycomb lattice.
Figure 3: (a) Left half shows the diffuse neutron scattering pattern in the ($h$, $k$, 0) plane for Cs3Fe2Cl9 measured at $T=6$ K on CORELLI. Data are symmetrized according to the $6/mmm$ Laue class. The right panel is the calculated diffuse neutron scattering pattern for the fitted $J_{1\textrm{-}5}$-$D_{z}$ model using the SCGA method assuming a reduced $T=5$ K to compensate for the underestimated critical correlations. Both the experimental and calculated data were integrated over l=[0.1,0.1]l=[-0.1,0.1] reciprocal lattice units, r.l.u. The solid-line hexagon indicates the boundary of the first Brillouin zone. Triangular-shaped lobes around the $K$-$(\frac{1}{3},\frac{1}{3},0)$ points are the spiral surface for a honeycomb-lattice model with a frustration ratio of $J_{2}^{\textrm{h}}/J_{1}^{\textrm{h}}=0.72$. (b) Similar experimental (left half) and calculated (right half) diffuse scattering patterns in the ($h$, $k$, 1) plane with an integration range of $l=[0.9,1.1]$ r.l.u. (c) Experimental (left half) and calculated (right half) diffuse scattering patterns with an integration range of l=[0.1,0.1]+[0.9,1.1]l=[-0.1,0.1]+[0.9,1.1] r.l.u. (d) Experimental (left half) and calculated (right half) diffuse scattering patterns in the $(h,h,l)$ plane with an integration width of 0.1 r.l.u. along (h,h,0)(h,-h,0). (e) Scattering intensity along $(h,h,0)$ integrated from the $(h,k,0)+(h,k,1)$ pattern in panel (c) with an integration width of 0.06 r.l.u. along (h,h,0)(h,-h,0) as outlined by the dashed-line rectangle in the left half of panel (c). (f) $l$-dependence of the scattering intensity integrated in the area of $[0.125,0.5]$ and [0.125,0.125][-0.125,0.125] r.l.u. along the $(h,h,0)$ and (h,h,0)(h,-h,0) directions, respectively. This area is outlined in the right half of panel (c) by a dashed-line rectangle. In panels (e) and (f), the red solid line shows the calculated results for the fitted $J_{1\textrm{-}5}$-$D_{z}$ model using the SCGA method.
@@
A special feature of the SSL in Cs3Fe2Cl9 is the visibility of the spiral surface, which arises from its unique codimension two. On the original honeycomb lattice with antiferromagnetic $J_{1}^{h}$, the spiral surface within the first Brillouin zone has a structure factor of zero due to the interference between the two sublattices Liu et al. (2016), leading to a diffuse pattern similar that shown in Fig. 3(a). However, as compared in Figs. 3(a) and (b), the visibility of the spiral surface in Cs3Fe2Cl9 is complementary between the $l=0$ and $l=1$ planes. This variance in visibility originates from the $l$-dependence of the phase factor, which modulates the interference between the two sublattices and thus reflects the higher codimension of the SSL Liu et al. (2016). By summing up the intensity in the planes of $l=0$ and 1, the complete spiral surface can be recovered in Fig. 3(c), upon which a line cut along the (h,h,0h,~h,~0) direction is shown in Fig. 3(e).
The scattering pattern in the (h,h,lh,~h,~l) plane presented in Fig. 3(d) reveals weak scattering intensity between the integer-$l$ planes. According to our calculations, the intensity of the interplanar scattering diminishes with decreasing temperatures, thus can be attributed to thermal excitations out of the ground state manifolds. This observation is further confirmed in Fig. 3(f) through the measured and calculated $l$-dependence of the integrated intensity within a rectangular area of $[0.125,0.5]$ r.l.u. and [0.125,0.125][-0.125,0.125] r.l.u. along the $(h,h,0)$ and (h,h,0)(h,-h,0) directions, respectively.
@@
Single crystal neutron diffraction experiments were performed on WAND2 Frontzek et al. (2018) at the High Flux Isotope Reactor (HFIR) at ORNL to clarify the LRO phases as a function of magnetic field up to 6 T sup . As summarized in Figs. 4(e)-4(h), four different types of diffraction patterns are observed in the ($h$, $k$, 0) plane among phases I-VI. Compared to that in phase I, the diffraction pattern of phase II in Fig. 4(f) exhibits additional weak magnetic Bragg peaks at $\bm{q}_{\textrm{II}}=(\frac{1}{4},\frac{1}{4},0)$, indicating the coexistence of minor $\uparrow\uparrow\uparrow\downarrow$ and major $\uparrow\downarrow\uparrow\downarrow$ magnetic domains, where $\uparrow$ ($\downarrow$) represents spins (antiparallel) parallel with $c$. Surprisingly, the diffraction patterns in phases III-V, among which the pattern for phase IV is shown in Fig. 4(g), are similar to each other sup : all exhibiting main reflections at $\bm{q}_{\textrm{III}}=\frac{5}{6}\times(\frac{1}{3},\frac{1}{3},0)$ close to the center of the edges of the triangular-shaped spiral surface as indicated by dashed lines, together with weaker secondary reflections at $2\bm{q}_{\textrm{III}}$. The existence of the latter often suggests a field-distorted spiral or SDW phase Gignoux et al. (1998); Stüßer et al. (2002).
As shown in Fig. 4(h) and also compared in Fig. 4(c) for the line cuts along the (h+1,h1,0)(h+1,h-1,0) direction, the $\bm{q}_{\textrm{III}}$ refelctions disappear in phase VI. Instead, new magnetic reflections belonging to $\bm{q}_{\textrm{VI}}=\frac{11}{9}\times(\frac{1}{3},\frac{1}{3},0)$ appear at the corners of the triangular-shaped spiral surface, which is exactly the position predicted by the ObD transition Mulder et al. (2010); Liu et al. (2020). Comparison of the line cuts along the (2h,4h,0)(-2h,4h,0) direction shown in Fig. 4(d) reveal a clear shift in the $\bm{q}_{\textrm{VI}}$ and $2\bm{q}_{\textrm{III}}$ positions, evidencing the distinct magnetic order in phase VI.
Using classical Monte Carlo simulations in a parallel tempering scheme as implemented in the SpinMC program Spi , we calculated the $H$-$T$ phase diagram for the fitted $J_{1\textrm{-}5}$-$D_{z}$ model. The calculated phase diagram, shown in Fig. 4(b), agrees well with the experimental results shown in Fig. 4(a), while some discrepancies existing in the transitional regime, including phases I’ and VI’ shadowed by dashed lines, may originate from the absence of quantum fluctuation in our calculations sup . The agreement between the experimental and calculated phase diagrams allows us to use the simulated magnetic structures as constraints in refining the magnetic orders against the neutron diffraction dataset collected in each phase sup .
Following this method, the magnetic orders in phases I-VI can be determined, and the representative structures in the neighboring bilayers are shown in Figs. 4(i)-(l). Phases III and IV are determined to be of an elongated spiral-type orders, and the transition between them can be attributed to the modulation of the $\bm{q}=0$ magnitude sup . As a contrast, both phases V and VI are revealed to be of the SDW orders with spins along the $c$ axis despite their different $\bm{q}$ vectors. Since $\bm{q}_{\textrm{VI}}$ in phase VI is the predicted propagation vector of the ObD theory Mulder et al. (2010); Liu et al. (2020), it is tantalizing to ascribe the transition from the SSL state to phase VI as entropy-driven. The proposed SDW character of phase VI, however, poses a challenge in accurately determining the free energy for states above $T_{N}$ under the current theoretical framework Bergman et al. (2007). Further studies will be required to clarify the character of the SSL-SDW phase transition.
Discussion. From the perspective of experimental realization, the introduction of codimension into SSLs is not a mere classification for convenience but rather has important implications Yao et al. (2021). On the original honeycomb lattice, the realization of SSL requires a relatively strong second-neighbor interactions, which is rarely satisfied in real materials Gao et al. (2022a). However, as demonstrated in our work for Cs3Fe2Cl9, the relatively strong intralayer couplings on the AB-stacked triangular lattice, which correspond to the second-neighbor interactions on an effective honeycomb lattice, naturally favor the emergence of SSLs. Besides the many compounds that are isostructural with Cs3Fe2Cl9 Brüning et al. (2021); Leuenberger et al. (1986a, b, c), there are a rich variety of candidate systems like RRZnPO Wan et al. (2024); Nientiedt and Jeitschko (1998) and $R_{3}$Ru4Al12 Hirschberger et al. (2019); Gao et al. (2019); Kolincio et al. (2021) where the SSL physics might be achieved.
Codimension-two SSLs also offer a plethora of novel physics that remain to be explored. One immediate question is the impacts of thermal and quantum fluctuations over the SSL state. Besides the thermally induced ObD transitions, there are theoretical discussions about the prospect of realization quantum spin liquid on the honeycomb lattice through SSLs Reuther et al. (2011); Niggemann et al. (2020). On the AB-stacked triangular lattice, this approach may still be valid, which resonates with the recent efforts of utilizing the further-neighbor interactions on a triangular lattice to achieve quantum spin liquids near the phase boundaries Gallegos et al. (2025). Another exciting direction to explore is the novel spin textures that may emerge from the codimension-two SSLs. On an AB-stacked triangular lattice, the inter- and intra-layer interactions may follow different symmetry constraints depending on the presence or location of the centrosymmetric points. This additional freedom in symmetry constraints can be highly beneficial for the realization of exotic spin textures like magnetic skyrmions Mohylna et al. (2022, 2025).